# An open-source heuristic to reboot 2D nesting research

### v2

@@ 1 Introduction
2D irregular cutting and packing (C&P) – or nesting – problems involve fitting a set of smaller items into larger containers, where the shape of the items and/or containers can be irregular (non-rectangular). Such optimization problems arise in a wide range of industries such as garment manufacturing, furniture production, shipbuilding, printing, laser-cutting, metalworking, woodworking, foam-cutting, leather crafting, and microchip design. Designing effective algorithms to solve these NP-hard problems (fowler1981optimal) has been an active area of research since the 1960s (art1966approach).
Given the real-world relevance of 2D irregular C&P problems and the ongoing shift in ethos toward open-source practices (unesco2021openscience), one would expect to find a vibrant ecosystem of publicly available implementations that respond to advances in the academic state of the art (SotA). However, despite more than a decade of stagnation – where there has been little to no improvement in achievable solution quality – the open-source landscape remains disconnected from academic progress and trails significantly behind in performance. We hypothesize that there are two interconnected factors primarily responsible for this situation.
@@
This paper has two interrelated goals: (i) bridge the gap between academic and open-source algorithms and (ii) lower the barrier to entry for future researchers. When addressing these goals, we will rely on the recently introduced Collision Detection Engine (CDE) for 2D irregular C&P problems (gardeyn2025decoupling), which can handle the geometric challenge. Our entire focus can therefore target the optimization challenge. On top of the CDE, we will design an open-source heuristic for the 2D irregular strip packing problem, arguably the most canonical 2D nesting problem. While we focus on strip packing, our broader ambition is to develop an algorithm that can generalize to other variants of 2D irregular C&P problems without much effort.
@@ 2 Problem definition
The objective of the 2DISPP is to determine a set of rigid transformations $\{t_{i}:i\in I\}$ such that, when each item $i$’s shape is transformed by its corresponding transformation $t_{i}$, all items are positioned entirely within the strip (Equation 1) and none collide with each other (Equation 2) for the smallest possible strip length $l$.
@@ 3 State of 2D irregular strip packing
Construction heuristics build solutions sequentially by placing one item at a time according to a predetermined ordering in addition to a placement rule. To improve solution quality, many methods embed a higher-level heuristic that explores different orderings and placement strategies. Representative work has been conducted by oliveira2000topos; gomes2002two; burke2006new; bennell2010beam; amaro2017parallel. These algorithms are generally straightforward to implement and reproducible provided that sufficient detail is included in the corresponding publication.
However, the achievable solution quality of construction heuristics lags significantly behind the SotA. We can attribute this to the nature of the 2DISPP, which does not lend itself well to such approaches for two main reasons. First, even minor changes to the ordering of items or their placement early in the construction process leads to unpredictable changes concerning final solution quality. This chaotic nature undermines the effectiveness of the higher-level heuristics.
Second, if one jumps ahead to the densely packed solutions in Figure 10, try to imagine the orderings and placement rules needed to position the items one by one into similar configurations. Maybe one could devise an effective placement strategy that works well for one of these instances, but it is hard to imagine a single set of rules that would work well for all.
@@
Most methods relying on iterative improvement begin with a complete solution and repeatedly apply local modifications – such as shifting or swapping items – in an attempt to find new solutions of higher quality. Early contributions include egeblad2007fast; umetani2009solving; leung2012extended. More recent advances by elkeran2013new; wang2017flexible; sato2019raster represent the SotA for the 2DISPP. The common theme across these methods is that they tolerate temporary collisions between items. The strip gets converted into a container with fixed dimensions and, every time a feasible solution is found, its length is reduced further. This shrinking leads to new collisions, which the aforementioned methods attempt to gradually resolve in order to reach a feasible solution for the shorter strip.
@@
Solid, robust and publicly available implementations to solve the 2DISPP are scarce. A notable exception is a tree-search heuristic by fontan. This heuristic has a unique approach to handling the geometric challenge that involves decomposing shapes into trapezoids. Despite representing a valuable open-source contribution, fontan’s method is unrelated to the iterative-improvement methods and cannot consistently compete with the academic SotA.
@@ 4 A sequence of feasibility problems
As indicated in Section 3, the best-performing algorithms for the 2DISPP temporarily relax the non-collision constraint (Equation 2) during their iterative search. These methods essentially convert the optimization problem into a sequence of feasibility problems. Each feasibility problem fixes strip length $l$ and the goal becomes to find a single solution that is feasible.
@@
The idea of exploring infeasible regions has existed for a long time (glover1989tabu) and is a strategy commonly used in the irregular C&P literature. However, the reasons for its effectiveness in nesting problems are rarely elaborated upon. We will therefore take this opportunity to offer some intuition for why this conversion works so well.
The 2DISPP is a highly constrained problem, with the non-collision constraint (Equation 2) being the most significant. This constraint is so restrictive that virtually any local modification, such as shifting or swapping the positions of items, would immediately render a solution infeasible. The high-quality solutions we seek are effectively rare oases in a vast desert of infeasibility. Algorithms that cannot venture outside the feasible region are prevented from making meaningful progress. The best-performing algorithms therefore temporarily relax this non-collision constraint, enabling them to traverse otherwise inaccessible routes towards high-quality solutions.
@@ 5 Detecting collisions
Recently, gardeyn2025decoupling introduced a Collision Detection Engine (CDE) to provide a fast and reliable way for 2D nesting algorithms to assess feasibility. This CDE builds on trigonometric principles and includes a comprehensive suite of algorithmic enhancements to overcome the traditional limitations of such trigonometric approaches. The CDE is implemented in an open-source library called jagua-rs [^1] https://github.com/JeroenGar/jagua-rs. jagua-rs supports irregular shapes for both items and containers, continuous rotation and translation, and easily extends to problem variants containing additional spatial constraints.
The optimization algorithm we will introduce in this paper will be built on top of jagua-rs: assessing feasibility through collision queries and issuing updates to reflect any changes in the solution. All interactions with jagua-rs throughout this paper can be abstracted into the following two functions:
@@ 6.1 Area of overlap
Figure 4(a) visualizes the output of this function for the pair of items from Figure 3, with the smaller item being positioned all around the larger one. The shapes and reference points of six example placements are visualized. The region where the two items do not collide is colored green. The region where they collide is shaded red, with the intensity of the red mapping to the output of the function described in Algorithm 2. To help visualize the intensity further, Figure 4(b) maps the value of this function to the height of a 3D surface.
@@ 6.2 Decaying penetration depth
$\delta^{\prime}=\begin{cases}\delta&\text{if }\delta\geq\varepsilon\\ \frac{\varepsilon^{2}}{-\delta+2\varepsilon}&\text{otherwise}\end{cases}$ (5)
@@ 7 Navigating a continuous search space
All that now remains is to design a search algorithm that is able to gradually resolve collisions. One of the most fundamental components of any local search algorithm is the move operation, enabling the algorithm to transition from one solution to a neighboring one. Algorithm 5 introduces the local search move our approach will employ. Its core idea – repositioning every colliding item, one by one, in a randomized order – is shared by umetani2009solving; sato2019raster; elkeran2013new.
@@
In 2D irregular C&P problems, items can be positioned in a continuous 2D space and sometimes also rotated in a continuous manner. This means it is far too time-consuming to evaluate a full neighborhood of all possible item positions. Previous approaches have handled this continuous search space in a variety of ways. sato2019raster discretize the search space into a (multi-resolution) grid. Meanwhile, approaches that rely on NFPs (umetani2009solving; elkeran2013new) narrow the search space down to a set of positions along the edges of the union of all NFPs. From this (still continuous) set of positions, only a few specific points such as the vertices of the combined NFP are typically considered. All of these approaches effectively restrict the search space to a finite set of positions.
@@
Algorithm 6 indicates how we search for a new position for a single item $i$. A sample is defined as an item and an accompanying transformation: $(i,t)$. We define a function to evaluate samples $f(i,t)\mapsto e$ which will, for now, remain abstract: it takes as input a sample and returns a single value $e$ indicating the quality of that sample. This function, along with the item, is passed to the sampling procedure, which returns $t^{*}$, the best position it found to reposition item $i$ to.
@@ 8.1 Guided local search
How to search for a new position was explained in Algorithm 6, but the function to evaluate samples was left abstract. A straightforward way to evaluate a sample would be to aggregate the severity of all collisions that would occur if the item were placed at the sampled position, akin to Algorithm 1. However, such a static evaluation function would cause the search to converge to a local optimum within a couple of iterations. Approaches proposed by umetani2009solving and sato2019raster avoid this issue by relying on Guided Local Search (GLS) (voudouris2010guided): a metaheuristic that dynamically modifies the evaluation function to escape local optima and move towards promising regions of the search space.
@@
5:    $m\leftarrow\texttt{M}_{l}+(\texttt{M}_{u}-\texttt{M}_{l})\cdot(e/e_{\text{max}})$
@@
7:    $m\leftarrow\texttt{M}_{d}$
@@
All item-pair weights are initialized to 1 and every time Algorithm 8 is called, each weight is updated by multiplying it by its corresponding factor $m$. If the pair of items is colliding, $m$ represents a linear mapping $\in\;\mathclose{]}\texttt{M}_{l},\texttt{M}_{u}\mathclose{]}$ of the severity of their collision $e$ relative to the most severe collision $e_{\text{max}}$. Item pairs with more severe collisions will thus have their weight rise faster than those with less severe collisions. If a pair is not in collision, their weight is multiplied by $\texttt{M}_{d}$ (<1<1) and eventually decays back to 1 over time.
@@ 8.2 Separation Procedure
Let us begin by defining the evaluation function for the entire solution:
@@
This function is the sum of the evaluation of every item pair (Algorithm 1). Whenever Equation 8 returns 0, all collisions have been resolved and a feasible solution has been reached.
@@
2: while $k<k_{\text{max}}$ and $e^{*}>0$ do
3:   restore to $s^{*}$} ; $s_{\text{init}}\leftarrow s^{*}$ ; $n\leftarrow 0$
4:   while $n<n_{\text{max}}$ and $e^{*}>0$ do $\triangleright$ an attempt
5:    move_items_multi $\triangleright$ Alg. 10
6:    update_weights $\triangleright$ Alg. 8
@@
8:    if $e<e^{*}$ then
9:      $s^{*}\leftarrow$ current solution ; $e^{*}\leftarrow e$
10:      $n\leftarrow 0$
11:   $k\leftarrow k+1$ $\triangleright$ add a strike
12:   if $s^{*}\neq s_{\text{init}}$ then
13:    $k\leftarrow 0$ $\triangleright$ reset strikes
14: return $s^{*}$
Algorithm 9 describes the complete procedure to solve the feasibility problem introduced in Section 4. The separation procedure starts from an infeasible situation with colliding items and attempts to resolve all of its collisions. At its core, the algorithm continually repositions items, updates the item-pair weights and replaces the incumbent solution $s^{*}$ whenever improvements are found with respect to Equation 8. Whenever feasibility is reached, the procedure stops and immediately returns the solution. Once a certain number of iterations without improvements $n_{\text{max}}$ is reached, we conclude one attempt. If the incumbent solution was never updated during this attempt, a strike is added. Conversely, if the attempt improved $s^{*}$, the strike counter $k$ is reset to 0. As long as the number of strikes is below a certain threshold $k_{\text{max}}$, the procedure restores the incumbent solution $s^{*}$ and starts a new attempt, continuing its quest for a feasible solution. Once the number of strikes exceeds $k_{\text{max}}$, the procedure halts the search and returns $s^{*}$, which will unfortunately still be infeasible.
@@ 8.3 Balancing exploration and exploitation
1: $s_{i}\leftarrow$ current solution ; $e^{*}\leftarrow\infty$
@@
5:   if $e<e^{*}$ then
6:    $s^{*}\leftarrow$ current solution ; $e^{*}\leftarrow e$
@@
Algorithm 10 essentially executes Algorithm 5 multiple times with different random orders of items. From these multiple neighboring solutions, the best one (Equation 8) is chosen. This significantly increases the chances of finding improvement while preserving the diversity of the search process. This enhanced version is employed in the separation procedure (Algorithm 9) to solve the feasibility problem. Although presented here sequentially, Algorithm 10 is parallelized across a pool of N_WORKERS threads in our implementation.

### v3

@@ 1 Introduction
2D irregular cutting and packing (C&P) – or nesting – problems involve fitting a set of smaller items into larger containers, where the shape of the items and/or containers can be irregular (non-rectangular). Such optimization problems arise in a wide range of industries such as garment manufacturing, furniture production, shipbuilding, printing, laser-cutting, metalworking, woodworking, foam-cutting, leather crafting, and microchip design. Designing effective algorithms to solve these NP-hard problems [fowler1981optimal] has been an active area of research since the 1960s [art1966approach].
Given the real-world relevance of 2D irregular C&P problems and the ongoing shift in ethos toward open-source practices [unesco2021openscience], one would expect to find a vibrant ecosystem of publicly available implementations that respond to advances in the academic state of the art. However, despite more than a decade of stagnation – where there has been little to no improvement in achievable solution quality – the open-source landscape remains disconnected from academic progress and trails significantly behind in performance. We hypothesize that there are two interconnected factors primarily responsible for this situation.
@@
This paper has two interrelated goals: (i) bridge the gap between academic and open-source algorithms and (ii) lower the barrier to entry for future researchers. When addressing these goals, we will rely on the recently introduced Collision Detection Engine (CDE) for 2D irregular C&P problems [gardeyn2025decoupling], which can handle the geometric challenge. Our entire focus can therefore target the optimization challenge. On top of the CDE, we will design an open-source heuristic for the 2D irregular strip packing problem, arguably the most canonical 2D nesting problem. While we focus on strip packing, our broader ambition is to develop an algorithm that can generalize to other variants of 2D irregular C&P problems without much effort.
Throughout this paper, we aim to present our method in an accessible manner, with an emphasis placed on high-level reasoning over exhaustive formal detail. While implementation details that would hinder readability are intentionally omitted, the accompanying open-source codebase serves as the formal specification of our method, with each algorithm explicitly referenced in the documentation. This creates two complementary resources: this paper provides intuition and understanding, while the accompanying codebase offers the formal specification for reproduction and implementation.
@@ 2 Problem definition
The objective of the 2DISPP is to determine a set of rigid transformations $\{t_{i}:i\in I\}$ such that, when each item $i$’s shape is transformed by its corresponding transformation $t_{i}$, all items are positioned entirely within the strip (Expression 1) and none collide with each other (Expression 2) for the smallest possible strip length $l$.
@@ 3 State of 2D irregular strip packing
Construction heuristics build solutions sequentially by placing one item at a time according to a predetermined ordering in addition to a placement rule. To improve solution quality, many methods embed a higher-level heuristic that explores different orderings and placement strategies. Representative work has been conducted by oliveira2000topos, gomes2002two, burke2006new, bennell2010beam, amaro2017parallel. These algorithms are generally straightforward to implement and reproducible provided that sufficient detail is included in the corresponding publication. However, the achievable solution quality of construction heuristics lags significantly behind the state of the art.
We can attribute this to the nature of the 2DISPP, which does not lend itself well to such approaches for two main reasons. First, even minor changes to the ordering of items or their placement early in the construction process leads to unpredictable changes concerning final solution quality. This chaotic nature undermines the effectiveness of the higher-level heuristics. Second, if one jumps ahead to the densely packed solutions in Figure 10, try to imagine the orderings and placement rules needed to position the items one by one into similar configurations. Maybe one could devise an effective placement strategy that works well for one of these instances, but it is hard to imagine a single set of rules that would work well for all.
@@
Most methods relying on iterative improvement begin with a complete solution and repeatedly apply local modifications – such as shifting or swapping items – in an attempt to find new solutions of higher quality. Early contributions include egeblad2007fast, umetani2009solving, leung2012extended. More recent advances by elkeran2013new, wang2017flexible, sato2019raster represent the state of the art for the 2DISPP. The common feature shared by these methods is that they tolerate temporary collisions between items. The strip gets converted into a container with fixed dimensions and, every time a feasible solution is found, its length is reduced further. This shrinking leads to new collisions, which the aforementioned methods attempt to gradually resolve in order to reach a feasible solution for the shorter strip.
@@
Solid, robust and publicly available implementations to solve the 2DISPP are scarce. A notable exception is a tree-search heuristic by fontan. This heuristic has a unique approach to handling the geometric challenge that involves decomposing shapes into trapezoids. Despite representing a valuable open-source contribution, fontan’s method is unrelated to the iterative-improvement methods and cannot consistently compete with the academic state of the art.
@@ 4 A sequence of feasibility problems
As indicated in Section 3, the best-performing algorithms for the 2DISPP temporarily relax the non-collision constraint (Expression 2) during their iterative search. These methods essentially convert the optimization problem into a sequence of feasibility problems. Each feasibility problem fixes strip length $l$ and the goal becomes to find a single solution that is feasible.
@@
The idea of exploring infeasible regions has existed for a long time [glover1989tabu] and is a strategy commonly used in the irregular C&P literature. However, the reasons for its effectiveness in nesting problems are rarely elaborated upon. We will therefore take this opportunity to offer some intuition for why this conversion works so well.
The 2DISPP is a highly constrained problem, with the non-collision constraint (Expression 2) being the most significant. This constraint is so restrictive that virtually any local modification, such as shifting or swapping the positions of items, would immediately render a solution infeasible. The high-quality solutions we seek are effectively rare oases in a vast desert of infeasibility. Algorithms that cannot venture outside the feasible region are prevented from making meaningful progress. The best-performing algorithms therefore temporarily relax this non-collision constraint, enabling them to traverse otherwise inaccessible routes towards high-quality solutions.
@@ 5 Detecting collisions
gardeyn2025decoupling introduced a Collision Detection Engine (CDE) to provide a fast and reliable way for 2D nesting algorithms to assess feasibility. This CDE builds on trigonometric principles and includes a comprehensive suite of algorithmic enhancements to overcome the traditional limitations of such trigonometric approaches. The CDE is implemented in an open-source library called jagua-rs [^1] https://github.com/JeroenGar/jagua-rs. jagua-rs supports irregular shapes for both items and containers, continuous rotation and translation, and easily extends to problem variants containing additional spatial constraints.
The optimization algorithm we will introduce in this paper, is built on top of jagua-rs: assessing feasibility through collision queries and issuing updates to reflect any changes in the solution. All interactions with jagua-rs throughout this paper can be abstracted into the following two functions:
@@ 6.1 Area of overlap
Figure 4(a) visualizes the output of this function for the pair of items from Figure 3, with the smaller item being positioned all around the larger one. The shapes and reference points of six example placements are shown. The region where the two items do not collide is colored green. The region where they collide is shaded red, with the intensity of the red mapping to the output of the function described in Algorithm 2. To help interpret the intensity further, Figure 4(b) maps the value of this function to the height of a 3D surface.
@@ 6.2 Decaying penetration depth
$\delta^{\prime}=\begin{cases}\delta&\text{if }\delta>\varepsilon\\ \frac{\varepsilon^{2}}{-\delta+2\varepsilon}&\text{otherwise}\end{cases}$ (5)
@@ 7 Navigating a continuous search space
All that now remains is to design a search algorithm that is able to gradually resolve collisions. One of the most fundamental components of any local search algorithm is the move operation, enabling the algorithm to transition from one solution to a neighboring one. Algorithm 5 introduces the local search move our approach will employ. Its core idea – repositioning every colliding item, one by one, in a randomized order – is shared by umetani2009solving, sato2019raster, elkeran2013new.
@@
In 2D irregular C&P problems, items can be positioned in a continuous 2D space and sometimes also rotated in a continuous manner. This means it is far too time-consuming to evaluate a full neighborhood of all possible item positions. Previous approaches have handled this continuous search space in a variety of ways. sato2019raster discretize the search space into a (multi-resolution) grid. Meanwhile, approaches that rely on NFPs [umetani2009solving, elkeran2013new] narrow the search space down to a set of positions along the edges of the union of all NFPs. From this (still continuous) set of positions, only a few specific points such as the vertices of the combined NFP are typically considered. All of these approaches effectively restrict the search space to a finite set of positions.
@@
Algorithm 6 indicates how we search for a new position (transformation) for an item $i$. A sample is defined as an item and an accompanying transformation: $(i,t)$. We define a function to evaluate samples $f(i,t)\mapsto e$ which will, for now, remain abstract: it takes as input a sample and returns a single value $e$ indicating the quality of that sample. This function, along with the item, is passed to the sampling procedure, which returns $t^{*}$, the best position it found to reposition item $i$ to.
@@ 8.1 Guided local search
How to search for a new position was explained in Algorithm 6, but the function to evaluate samples was left abstract. A straightforward way to evaluate a sample would be to aggregate the severity of all collisions that would occur if the item were placed at the sampled position, akin to Algorithm 1. However, such a static evaluation function would cause the search to converge to a local optimum within a couple of iterations. Approaches proposed by umetani2009solving and sato2019raster avoid this issue by relying on Guided Local Search (GLS) [voudouris2010guided]: a metaheuristic that dynamically modifies the evaluation function to escape local optima and move towards promising regions of the search space.
@@
5:    $m\leftarrow\texttt{M}_{l}+(\texttt{M}_{u}-\texttt{M}_{l})\cdot(e/e_{\text{max}})$ $\triangleright$ Table 1
@@
7:    $m\leftarrow\texttt{M}_{d}$ $\triangleright$ Table 1
@@
All item-pair weights are initialized to 1 and every time Algorithm 8 is called, each weight is updated by multiplying it by its corresponding factor $m$. If the pair of items is colliding, $m$ represents a linear mapping $\in\;\mathclose{]}\texttt{M}_{l},\texttt{M}_{u}\mathclose{]}$ of the severity of their collision $e$ relative to the most severe collision $e_{\text{max}}$. Item pairs with more severe collisions will thus have their weight rise faster than those with less severe collisions. If a pair is not in collision, their weight is multiplied by $\texttt{M}_{d}$ (<1<1) and eventually decays back to 1 over time. Parameters $\texttt{M}_{l}$, $\texttt{M}_{u}$ and $\texttt{M}_{d}$ are defined in Table 1.
@@ 8.2 Separation Procedure
Let us begin by defining the global objective function to evaluate an entire solution:
@@
This function is an expression of the total severity of all collisions in the solution, computed as the sum of the evaluation of every item pair (Algorithm 1). Whenever $z=0$, all collisions have been resolved and a feasible solution has been reached.
@@
2: $w_{ab}\leftarrow 1$ for all item pairs $a,b\in I$ $\triangleright$ GLS weights
3: while $k<k_{\text{max}}$ and $z^{*}>0$ do
4:   restore to $s^{*}$} ; $s_{\text{init}}\leftarrow s^{*}$ ; $n\leftarrow 0$
5:   while $n<n_{\text{max}}$ and $z^{*}>0$ do $\triangleright$ an attempt
6:    move_items_multi $\triangleright$ Alg. 10
7:    update_weights $\triangleright$ Alg. 8
@@
9:    if $z<z^{*}$ then $\triangleright$ improvement found
10:      $s^{*}\leftarrow$ current solution ; $z^{*}\leftarrow z$
11:      $n\leftarrow 0$
12:   $k\leftarrow k+1$ $\triangleright$ add a strike
13:   if $s^{*}\neq s_{\text{init}}$ then
14:    $k\leftarrow 0$ $\triangleright$ reset strikes
15: return $s^{*}$
Algorithm 9 describes the complete procedure to solve the feasibility problem introduced in Section 4. The separation procedure represents a local search heuristic that starts from an infeasible situation with colliding items and attempts to resolve all of its collisions. At its core, the algorithm continually repositions items, updates the item-pair weights and replaces the incumbent solution $s^{*}$ whenever improvements are found with respect the objective function in Expression 8. Whenever feasibility is reached, the procedure stops and immediately returns the solution. Once a certain number of iterations without improvements $n_{\text{max}}$ is reached, we conclude one attempt. If the incumbent solution was never updated during this attempt, a strike is added. Conversely, if the attempt improved $s^{*}$, the strike counter $k$ is reset to 0. As long as the number of strikes is below a certain threshold $k_{\text{max}}$, the procedure restores the incumbent solution $s^{*}$ and starts a new attempt, continuing its quest for a feasible solution. Once the number of strikes exceeds $k_{\text{max}}$, the procedure halts the search and returns $s^{*}$, which will unfortunately still be infeasible.
@@ 8.3 Balancing exploration and exploitation
1: $s_{i}\leftarrow$ current solution ; $z^{*}\leftarrow\infty$
@@
5:   if $z<z^{*}$ then
6:    $s^{*}\leftarrow$ current solution ; $z^{*}\leftarrow z$
@@
Algorithm 10 essentially executes Algorithm 5 multiple times with different random orders of items. From these multiple neighboring solutions, the best one (Expression 8) is chosen. This significantly increases the chances of finding improvement while preserving the diversity of the search process. This enhanced version is employed in the separation procedure (Algorithm 9) to solve the feasibility problem. Although presented here sequentially, Algorithm 10 is parallelized across a pool of N_WORKERS threads in our implementation.